# Statistical Inferences and Predictions for Areal Data and Spatial Data Fusion with Hausdorff–Gaussian Processes

### v2

@@ 5.2 Data fusion: Air pollution in Ventura and Los Angeles
Our second application analyzes fused fine particulate matter (PM2.5) data in Ventura and Los Angeles counties, California, USA. The dataset combines point-referenced direct measurements (EPA, 2018) from 1999 to present with areal satellite-derived estimates (Van Donkelaar et al., 2015), averaged from 2010 to 2012 (Figure 5). To align temporally, we averaged the direct measurements over the same period. The study region comprises 19 measurement stations and 184 grid boxes (approximately 101.95 km2 each), excluding islands. Our primary goal is to predict PM2.5 concentrations at a finer scale (approximately 45 km2) using both data sources. Secondarily, we aim to estimate average PM2.5 concentration and its spatial dependence across both counties.
Building on previous concepts, we introduce the notation for this application. The response variable is the recorded PM2.5, noted $Y(\mathbf{s}_{i})$ at the spatial unit $\mathbf{s}_{i}$, which is either a point or a grid box. We again resort to the GLMM model from Section 3.1. In this case, there are no covariates, the likelihood is Normal with variance $\tau^{2}$, and the link function is the identity. A heteroscedastic HGP prior is placed on the random effect allowing for differing variances between point-referenced and areal data through the SD function from Equation (2) with $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)\}$. For interpretability, we present results for $\sigma=\exp\{\alpha_{0}\}$ and $\sigma_{a}=\exp\{\alpha_{0}+\alpha_{1}\}$ instead of $\alpha_{0}$ and $\alpha_{1}$. The model was fit with $\nu$ fixed at different values, and we selected the model with best goodness-of-fit ($\nu=0.8$). For more details, see Table LABEL:tab:sm:gof_la in the appendix. The priors for model parameters are as in Section 4.2. The MCMC algorithm initialization is described in Section 3.2. We used four parallel chains, each comprising a warm-up period of 2000 samples and an additional 2000 samples for inference. All model parameters exhibited split-$\hat{R}$ values below, indicating good convergence. This is further supported by the trace plots presented in Section LABEL:sec:sm:df_convergence of the appendix.
For comparison, we also fit an AGP random effect model, with priors for the model parameters as described in Section 4.2. Two meshes of different resolutions were considered; see Figure LABEL:fig:sm:meshes in the appendix. The sparse- and fine-mesh AGP models are denoted AGP1 and AGP2, respectively.
@@
Table 3 also presents the out-of-sample prediction performance of the three models, assessed via 10-fold cross-validation. Using four metrics from our prior simulation study (RMSP, CPP, prediction interval width, IS), we find the HGP offers the most reliable point predictions, evidenced by its lowest RMSP. The HGP also maintains nominal coverage for its prediction intervals. While the AGP2 achieves similar coverage, its prediction intervals are considerably wider, resulting in inferior interval scores. Further analysis by data type is detailed in Table LABEL:tab:sm:out-pred_df of the appendix, revealing the unreliability of competing methods for point-referenced data, where frequentist coverage can be as low as 3%.
The drastic differences in the AGP results for different mesh resolutions are striking. The credible interval width for the intercept varies significantly between the two models. Crucially, the practical range reveals a key divergence: AGP1 indicates moderate spatial dependence, while AGP2 suggests weak dependence, likely explaining the observed overfitting in the latter. This weak dependence may imply AGP2 assigns nearly independent random intercepts to each data point, with variances inversely proportional to the sample unit’s area. AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (Fuentes and Raftery, 2005; Liu et al., 2011; Wang and Furrer, 2021).
@@ 6 Discussion
The proposed framework unifies the modeling of spatial data, extending geostatistical models to include areal and fused data within the same setting. We established conditions under which the PEC function of the HGP is positive definite, ensuring its validity. As a prior for spatial random effects in a GLMM, the HGP proves to be competitive to specialized models for diverse spatial data types. Our simulation studies indicate that the HGP is particularly advantageous for areal data with units of varying sizes and shapes, as confirmed by our disease mapping application. Unlike its specialized counterparts, the HGP also provides a directly interpretable spatial parameter, offering insights into the practical range of spatial dependence. In data fusion contexts, the HGP delivers more reliable interval predictions compared to the sparse AGP, and it provided similar results to the fine AGP in parameter estimation and spatial range. Additionally, in the air pollution data application, the proposed model consistently outperformed the AGP regardless of the resolution of its mesh. This success is even more remarkable considering the HGP bypasses the subjective step of defining a grid or mesh, a crucial but problematic requirement for the AGP. In addition, independent of the spatial data type, it continues to have direct and interpretable parameters.
@@
The HGP framework offers numerous opportunities for development. Further researching its theoretical properties will strengthen the foundation for its applications. A crucial focus should be developing correlation functions designed specifically for the HGP, particularly those suited to scenarios where spatial data is observed on a sphere. Approximations inspired by methods like the Nearest Neighbor Gaussian Process (Datta et al., 2016; Quiroz et al., 2023) or reduced rank approximations (Solin and Särkkä, 2020) could significantly broaden the applicability of the HGP to large-scale datasets. These advancements hold the potential to establish the HGP as an even more powerful and versatile tool for spatial data analysis.

### v3

@@ 5.2 Data fusion: Air pollution in Ventura and Los Angeles
Our second application analyzes fused fine particulate matter (PM2.5) data in Ventura and Los Angeles counties, California, USA. The dataset combines point-referenced direct measurements (U.S. Environmental Protection Agency Clean Air Markets Division,, 2018) from 1999 to present with areal satellite-derived estimates (Van Donkelaar et al.,, 2015), averaged from 2010 to 2012 (Figure 6). To align temporally, we averaged the direct measurements over the same period. The study region comprises 19 measurement stations and 184 grid boxes (approximately 101.95 km2 each), excluding islands. Our primary goal is to predict PM2.5 concentrations at a finer scale (approximately 45 km2) using both data sources. Secondarily, we aim to estimate average PM2.5 concentration and its spatial dependence across both counties.
Building on previous concepts, we introduce the notation for this application. The response variable is the recorded PM2.5, noted $Y(\mathbf{s}_{i})$ at the spatial unit $\mathbf{s}_{i}$, which is either a point or a grid box. We again resort to the GLMM model from Section 3.1. In this case, there are no covariates, the likelihood is Normal with variance $\tau^{2}$, and the link function is the identity. A heteroscedastic HGP prior is placed on the random effect allowing for differing variances between point-referenced and areal data through the SD function from Equation (2) with $w(\mathbf{s})=\mathbbm{1}(\mathcal{A}(\mathbf{s})>0)$. For interpretability, we present results for $\sigma=\exp\{\alpha_{0}\}$ and $\sigma_{a}=\exp\{\alpha_{0}+\alpha_{1}\}$ instead of $\alpha_{0}$ and $\alpha_{1}$. The model was fit with $\nu$ fixed at different values, and we selected the model with best goodness-of-fit ($\nu=0.8$). For more details, see Table S.5 in the Appendix Material. The priors for model parameters are as in Section 4.2. The MCMC algorithm initialization is described in Section 3.2. We used four parallel chains, each comprising a warm-up period of 2000 samples and an additional 2000 samples for inference. All model parameters exhibited split-$\hat{R}$ values below, indicating good convergence. This is further supported by the trace plots presented in Section S.6.1 of the Appendix Material.
For comparison, we also fit an AGP random effect model, with priors for the model parameters as described in Section 4.2. Two meshes of different resolutions were considered; see Figure S.17 in the Appendix Material. The sparse- and fine-mesh AGP models are denoted AGP1 and AGP2, respectively.
@@
Table 3 also presents the out-of-sample prediction performance of the three models, assessed via 10-fold cross-validation. Using four metrics from our prior simulation study (RMSP, CPP, prediction interval width, IS), we find the HGP offers the most reliable point predictions, evidenced by its lowest RMSP. The HGP also maintains nominal coverage for its prediction intervals. While the AGP2 achieves similar coverage, its prediction intervals are considerably wider, resulting in inferior interval scores. Table S.6 in the Appendix Material provides a breakdown by data type, highlighting the poor performance of competing methods for point-referenced data. The PI frequentist coverage, for example, can be as low as 3%.
The drastic differences in the AGP results for different mesh resolutions are striking. The credible interval width for the intercept varies significantly between the two models. Crucially, the practical range reveals a key divergence: AGP1 indicates moderate spatial dependence, while AGP2 suggests weak dependence, likely explaining the observed overfitting in the latter. This weak dependence may imply AGP2 assigns nearly independent random intercepts to each data point, with variances inversely proportional to the sample unit’s area. AGP mesh resolution is typically practitioner-informed, and results are highly sensitive to this tuning parameter, which lacks a consensus on “optimal resolution” (Fuentes and Raftery,, 2005; Liu et al.,, 2011; Wang and Furrer,, 2021).
@@ 6 Discussion
The proposed framework unifies the modeling of spatial data, extending geostatistical models to include areal and fused data within the same setting. As a prior for spatial random effects in a GLMM, the HGP proves to be competitive to specialized models for diverse spatial data types. Our simulation studies indicate that the HGP is particularly advantageous for areal data with units of varying sizes and shapes, as confirmed by our disease mapping application. Unlike its specialized counterparts, the HGP also provides a directly interpretable spatial parameter, offering insights into the practical range of spatial dependence. In data fusion contexts, the HGP delivers more reliable interval predictions compared to the sparse AGP, and it provided similar results to the fine AGP in parameter estimation and spatial range. Additionally, in the air pollution data application, the proposed model consistently outperformed the AGP regardless of the resolution of its mesh. This success is even more remarkable considering the HGP bypasses the subjective step of defining a grid or mesh, a crucial but problematic requirement for the AGP. In addition, independent of the spatial data type, it continues to have direct and interpretable parameters.
@@
The HGP framework offers numerous opportunities for development. Further researching its theoretical properties will strengthen the foundation for its applications. A crucial focus should be developing correlation functions designed specifically for the HGP, particularly those suited to scenarios where spatial data is observed on a sphere. Approximations inspired by methods like the Nearest Neighbor Gaussian Process (Datta et al.,, 2016; Quiroz et al.,, 2023) or reduced rank approximations (Solin and Särkkä,, 2020) could significantly broaden the applicability of the HGP to large-scale datasets. These advancements hold the potential to establish the HGP as an even more powerful and versatile tool for spatial data analysis.